# Configuration spaces of circles in the plane

### v1

@@ 5.1. The unlabeled case
$0\to\pi_{1}(F,\varkappa_{T})\to\pi_{1}(\operatorname{Conf}_{T}(S^{1},\mathbb{R}^{2}),\varkappa_{T})\to B_{m}^{\Pi}\to 0\text{.}$
Recall that the connected component of $F$ containing $\varkappa_{T}$ is homotopy equivalent to $\operatorname{Conf}_{T(v_{1})}(S^{1},\mathbb{R}^{2})\times\cdots\times\operatorname{Conf}_{T(v_{m})}(S^{1},\mathbb{R}^{2})$. Under this homotopy equivalence, the basepoint $\varkappa_{T}$ gets identified with $(\varkappa_{T(v_{1})},\dots,\varkappa_{T(v_{m})})$ thanks to the construction in Definition 3.3. Thus, by induction, our short exact sequence becomes
$0\to\operatorname{BAut}(T(v_{1}))\times\cdots\times\operatorname{BAut}(T(v_{m}))\to\pi_{1}(\operatorname{Conf}_{T}(S^{1},\mathbb{R}^{2}),\varkappa_{T})\to B_{m}^{\Pi}\to 0\text{.}$
The map $\pi_{1}(\operatorname{Conf}_{T}(S^{1},\mathbb{R}^{2}),\varkappa_{T})\to B_{m}^{\Pi}$ clearly splits, so by now we know that our group of interest is isomorphic to
@@
which is exactly the form for $\operatorname{BAut}(T)$ in Definition 4.4. The only thing left to prove is that the action of $B_{m}^{\Pi}$ on $\operatorname{BAut}(T(v_{1}))\times\cdots\times\operatorname{BAut}(T(v_{m}))$ here is the same as the action of $B_{m}^{\Pi}$ on $\operatorname{BAut}(T(v_{1}))\times\cdots\times\operatorname{BAut}(T(v_{m}))$ in the definition of $\operatorname{BAut}(T)$, i.e., it is by permuting entries. But this is exactly what is ensured by the above splitting, where the non-top circles “go along for the ride” as the top circles move. ∎
@@ 5.2. The labeled case
Let $T$ be a labeled tree with $n$ non-root vertices. Then $\operatorname{LConf}_{T}(S^{1},\mathbb{R}^{2})$ is aspherical, and its fundamental group is isomorphic to $\operatorname{PBAut}(T)$.
@@
Since $\operatorname{LConf}_{T}(S^{1},\mathbb{R}^{2})$ is a covering space of $\operatorname{Conf}_{T}(S^{1},\mathbb{R}^{2})$, which is aspherical by Theorem A, we know that $\operatorname{LConf}_{T}(S^{1},\mathbb{R}^{2})$ is also aspherical.
As for the fundamental group, by covering theory together with the correspondence in the unlabeled case from Theorem A, we have that it is isomorphic to the kernel of $\pi\colon\operatorname{BAut}(T)\to\operatorname{Aut}(T)$, which is $\operatorname{PBAut}(T)$. ∎

### v2

@@ 5.1. The unlabeled case
$0\to\pi_{1}(F,\underline{\kappa}_{T})\to\pi_{1}(\operatorname{UConf}_{T}(S^{1},\mathbb{R}^{2}),\underline{\kappa}_{T})\to B_{m}^{\Pi}\to 0\text{.}$
Again by Lemma 5.2, the connected component of $F$ containing $\underline{\kappa}_{T}$ is homotopy equivalent to $\operatorname{UConf}_{T(v_{1})}(S^{1},\mathbb{R}^{2})\times\cdots\times\operatorname{UConf}_{T(v_{m})}(S^{1},\mathbb{R}^{2})$. Moreover, under this homotopy equivalence the basepoint $\underline{\kappa}_{T}$ is identified with $(\underline{\kappa}_{T(v_{1})},\dots,\underline{\kappa}_{T(v_{m})})$ thanks to Construction 3.3. Thus, by induction, our short exact sequence becomes
$0\to\operatorname{bAut}(T(v_{1}))\times\cdots\times\operatorname{bAut}(T(v_{m}))\to\pi_{1}(\operatorname{UConf}_{T}(S^{1},\mathbb{R}^{2}),\underline{\kappa}_{T})\to B_{m}^{\Pi}\to 0\text{.}$
The map $\pi_{1}(\operatorname{UConf}_{T}(S^{1},\mathbb{R}^{2}),\underline{\kappa}_{T})\to B_{m}^{\Pi}$ splits, since sending the element of $B_{m}^{\Pi}$ represented by a path $\gamma(t)$ to the element of $\pi_{1}(\operatorname{UConf}_{T}(S^{1},\mathbb{R}^{2}),\underline{\kappa}_{T})$ represented by the path $\widetilde{\gamma}(t)$ constructed above gives a well defined group homomorphism. By now we know that our group of interest is isomorphic to
@@
which is exactly the form for $\operatorname{bAut}(T)$ in Definition 4.4. The only thing left to prove is that the action of $B_{m}^{\Pi}$ on $\operatorname{bAut}(T(v_{1}))\times\cdots\times\operatorname{bAut}(T(v_{m}))$ here is the same as the action of $B_{m}^{\Pi}$ on $\operatorname{bAut}(T(v_{1}))\times\cdots\times\operatorname{bAut}(T(v_{m}))$ in the definition of $\operatorname{bAut}(T)$, i.e., it is by permuting entries. Indeed, as explained above, the element of $B_{m}^{\Pi}$ represented by $\gamma(t)$ acts via the induced bijection $\sigma_{\gamma}\in S_{m}$. ∎
@@ 5.2. The labeled case
Let $T$ be a labeled tree with $n$ non-root vertices. Then $\operatorname{Conf}_{T}(S^{1},\mathbb{R}^{2})$ is aspherical, and its fundamental group is isomorphic to $\operatorname{pbAut}(T)$.
@@
Write $\overline{T}$ for the underlying tree of $T$. Since $\operatorname{Conf}_{T}(S^{1},\mathbb{R}^{2})$ is a covering space of $\operatorname{UConf}_{\overline{T}}(S^{1},\mathbb{R}^{2})$, which is aspherical by Theorem A, we know that $\operatorname{Conf}_{T}(S^{1},\mathbb{R}^{2})$ is also aspherical.
In order to understand the fundamental group we need some more information about this covering. First we have the covering $\operatorname{Conf}_{n}(S^{1},\mathbb{R}^{2})\to\operatorname{UConf}_{n}(S^{1},\mathbb{R}^{2})$, which sends $(C_{1},\dots,C_{n})$ to $\{C_{1},\dots,C_{n}\}$, so this has deck group $S_{n}$, acting by permuting the $C_{i}$. Restricting to the covering map $\operatorname{Conf}_{T}(S^{1},\mathbb{R}^{2})\to\operatorname{UConf}_{\overline{T}}(S^{1},\mathbb{R}^{2})$, the deck group is the subgroup of $S_{n}$ that, in permuting the $C_{i}$, preserves the nesting relation. This subgroup of $S_{n}$ is precisely $\operatorname{Aut}(\overline{T})$. This action is transitive on the fiber in $\operatorname{Conf}_{T}(S^{1},\mathbb{R}^{2})$ of any given $\{C_{1},\dots,C_{n}\}$, hence the covering is a regular covering. By [Hat02][Proposition 1.39] we get that $\pi_{1}(\operatorname{Conf}_{T}(S^{1},\mathbb{R}^{2}))$ is isomorphic to the kernel of $\pi_{1}(\operatorname{UConf}_{\overline{T}}(S^{1},\mathbb{R}^{2}))\to\operatorname{Aut}(\overline{T})$. The isomorphism $\pi_{1}(\operatorname{UConf}_{\overline{T}}(S^{1},\mathbb{R}^{2}))\cong\operatorname{bAut}(\overline{T})$ constructed in the proof of Theorem A is compatible with the respective maps to $\operatorname{Aut}(\overline{T})$, and so we get that $\pi_{1}(\operatorname{Conf}_{T}(S^{1},\mathbb{R}^{2}))$ is isomorphic to the kernel of $\pi\colon\operatorname{bAut}(\overline{T})\to\operatorname{Aut}(\overline{T})$, which is $\operatorname{pbAut}(T)$. ∎